OpenStudy (anonymous):

the product of 2 positive intergers is 90 find the 2 intergers if the larger is 3 more than twice the smaller

OpenStudy (mertsj):

x= the smaller 2x+3= the larger \[x(2x+3)=90\]

OpenStudy (mertsj):

\[2x^2+3x-90=0\]

OpenStudy (mertsj):

\[(2x+15)(x-6)=0\]

OpenStudy (mertsj):

\[2x+15=0\] \[x-6=0\]

OpenStudy (mertsj):

Discard 2x+15=0 because -15/2 is not positive and is not an integer.

OpenStudy (mertsj):

\[x=6\]

OpenStudy (mertsj):

\[2x+3=15\]

OpenStudy (mertsj):

So the integers and 6 and 15

OpenStudy (anonymous):

it was cut off the work of the intergers